OpenStudy (anonymous):

A small bag of sand is released from an ascending hot-air balloon whose upward constant velocity is v0 = 1.95 m/s. Knowing that at the time of the release the balloon was 57.8 m above the ground, determine the time, τ, it takes the bag to reach the ground from the moment of its release.

OpenStudy (anonymous):

@Shalante

OpenStudy (matt101):

When the bag leaves the balloon, the only force acting on it is gravity. We know the following information: initial velocity (vi), distance (d), acceleration (a). We want to calculate time (t). The equation that uses all these variables is: \[d=v_it+\frac{1}{2}at^2\]

OpenStudy (anonymous):

\[57.8m =(1.95 m/s)(t) +\frac{ 1 }{ 2 }(-9.81 m/s ^{2})(t ^{2})\]

OpenStudy (anonymous):

But how do I solve it exactly? As in getting the time all to one side and such?

OpenStudy (anonymous):

@Data_LG2 @whpalmer4 @matt101

OpenStudy (anonymous):

do you know how to solve a quadratic formula? you have to use the same method for this problem

OpenStudy (anonymous):

It's been awhile since I've done that.

OpenStudy (anonymous):

\(\sf 57.8m =(1.95 m/s)(t) +\frac{ 1 }{ 2 }(-9.81 m/s ^{2})(t ^{2})\\ -4.91t ^{2} + 1.95t- 57.8 =0 \) Use the quadratic formula to solve for t: \(\sf \Large t=\frac{ -b \pm \sqrt{b^2-4ac} }{2a}\) where a= -4.9, b=1.95, c= -57.8 you just have to plug these value in the formula to calculate for t. you must get two values: one negative and one positive value. But in this case, we must ignore the negative value since there is no negative time...

OpenStudy (anonymous):

I got 4.78s

OpenStudy (anonymous):

Is that the wrong answer?

OpenStudy (anonymous):

I think I made a mistake. Displacement is towards down so it should be -57.8m hence, the equation must be: \(\sf -57.8m =(1.95 m/s)(t) +\frac{ 1 }{ 2 }(-9.81 m/s ^{2})(t ^{2})\\ -4.91t ^{2} + 1.95t +57.8 =0\) you will get an imaginary number if you didn't follow the right sign convention. You already used -9.8 m/s^2, which means your that your upward direction is your positive.

OpenStudy (anonymous):

So am I needing to switch the -57.8 to a positive and plug it back in into the quadratic equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get -77.82 that way.

OpenStudy (anonymous):

no.. can you show me your solution?

OpenStudy (anonymous):

\[-1.95+\sqrt{(1.95^{2}-4(-4.9)(57.8)}/2(-4.9)\]

OpenStudy (anonymous):

You divide the entire thing by 2 (-4.9)

OpenStudy (anonymous):

remember that a quadratic can have two solutions, as I mentioned above, you are expected to get a negative and a positive value. Notice that in the quadratic formula that it has "\(\sf \color{red} \pm\)" symbol which means to add and subtract. You just did the "+" part.. you also have to do this \(\sf \large \frac{-1.95 \color{red}{-}\sqrt{(1.95^{2})-4(-4.9)(57.8)}\ }{2(-4.9)}\)

OpenStudy (anonymous):

87.38s

OpenStudy (anonymous):

I got 3.64s. Maybe you made some mistakes on your calculation, try checking it again and let me know if you already got it :)

OpenStudy (anonymous):

I keep getting the 87.38

OpenStudy (anonymous):

okay, let's do it step by step Inside the radical first: 1.95^2 = 3.8025 then -4(-4.9)(57.8)= 1132.88 add those two together 1136.6825 if you get the square root of it, what will you get?

OpenStudy (anonymous):

I got it finally!

OpenStudy (anonymous):

3.6392

OpenStudy (anonymous):

yay \.^_^./ just don't forget the unit.

OpenStudy (anonymous):

Right. Thanks for helping on this. Do you have time to help with another?

OpenStudy (anonymous):

not sure though, i'm doing my physics lab report. Just post it, maybe other users who are good with physics can help you too :)

OpenStudy (anonymous):

Alright. Thanks!